The molefraction of methanol in its 4.5 molal aqueous solution is ___________

Grade:12th pass

4 Answers

Mohd Mujtaba
6 years ago
Molality =no of moles of solute/mass of solvent (in kg) .here Molality =4.5 So, it means 4.5 moles of solute present in 1kg solvent. Mole of solvent=1000/18=55.5 So Mole fraction of solute=4.5/4.5+ 55.5=4.5/60=3/40=""0.075."" Thank☺☺☺☺
sneha
6 years ago
Thank you soooooo much............u solved my confusion. ........what is ur name ? Where r u frm? I am frm kerala. Thanks a lot 😊😊😊😊😊😆😆😃😀😃🍀🌼☺☺☺☺
Mohd Mujtaba
6 years ago
I am happy that my answer helped you.thank a lot.my name is mujtaba and I from UP(allahabad)☺☺😊😊😊😊😊askiitian can help you a lot in your query
sneha
6 years ago
oh!!! yup askiitian is very helpful ...and nice to meet u...r u in 12??or  r u practising for entrance???iam repeating neet entrance...